OpenStudy (anonymous):

Find the midpoint -7,-6 -2,10

OpenStudy (anonymous):

10 - - 6 = 16 16/2 = 8 -2 - - 7 = 5 5/2 = 5/2 answer = 5/2 , 8 i think

sam (.sam.):

\[\huge M=(\frac{x_{1}+x_{2}}{2},\frac{y_{1}+y_{2}}{2})\]

sam (.sam.):

M=(-3,-8)